OpenStudy (anonymous):

A car loan requires equal payments of $304 per month for four years. A. What is the total amount paid for four years? B. suppose a down payment of $4,200 was made in addition to the monthly payments. What was the total cost of the car?

OpenStudy (anonymous):

12*4=48 48*304=14592=A 14592-4200=10392=B